St. Kilda Enterprises produces parts for the electronics industry. The production manager and cost analyst reviewed the accounts for the previous month and have provided an estimated breakdown of the fixed and variable portions of manufacturing overhead.

Fixed Variable Total
Indirect materials $ 5,000 $ 10,000 $ 15,000
Indirect labor 3,000 17,000 20,000
Supervision 11,000 4,000 15,000
Depreciation 38,000 6,000 44,000
Maintenance 18,000 23,000 41,000
Total $ 75,000 $ 60,000 $ 135,000

Direct materials for the month amounted to $107,500. Direct labor for the month was $202,500. During the month, 12,500 units were produced.


Required:

a. No changes are expected in these cost relations next month. The firm has budgeted production of 16,250 units. Provide an estimate for total production cost for next month.
b. Determine the cost per unit of production for the previous month and the next month.
Answer #1

Variable cost changes with the change in output while total fixed cost remains same within the relevant range

Manufacturing overhead for next month = 75,000 + 60,000*16,250/12,500

= $153,000

Direct material cost = 107,500*16,250/12,500 = $139,750

Direct Labor = 202,500*16,250/12,500 = $263,250

Total production cost for next month = $556,000

b.Cost per unit:

Previous month = 445,000/12,500 = $35.6

Next month = 556,000/16,250 = $34.22